OpenStudy (anonymous):

Let f(x)=x^2x Use logarithmic differentiation to determine the derivative ƒ'(x) ?

OpenStudy (experimentx):

lake log on both sides, log y = 2x log x use implicit differentiation your answer will be (2logx+2)x^2x

OpenStudy (anonymous):

THANKS A LOT! GNITE!!

OpenStudy (anonymous):

DO I HAVE TO TAKE THE DERIVative also?

OpenStudy (experimentx):

yes ,,, dy/dx*1/y = (2logx+2) or dy/dx = y (2logx+2) y = x^2x or dy/dx = x^2x(2logx+2)

OpenStudy (anonymous):

hmmm

OpenStudy (anonymous):

x^2 x (2 log(x)+2) ??

OpenStudy (experimentx):

x^( 2x) * (2 log(x)+2)

OpenStudy (anonymous):

thanks!!

OpenStudy (anonymous):

there is a formula for such types of differentiation also,it goes as d/dx(u^v)= u^v*d/dx(v*log u) apply it here and see what you get???